Not all visitors to a certain company's website are customers. In fact, the website administrator estimates that about 12% of all visitors to the website are looking for other websites. Assuming that this estimate is correct, find the probability that, in a random sample of 5 visitors to the website, exactly 4 actually are looking for the website.
Round your response to at least three decimal places.

Answer by stanbon(75887)

Not all visitors to a certain company's website are customers. In fact, the website administrator estimates that about 12% of all visitors to the website are looking for other websites. Assuming that this estimate is correct, find the probability that, in a random sample of 5 visitors to the website, exactly 4 actually are looking for the website.
Round your response to at least three decimal places.
-----------------------------
It's a binomial problem with n = 5, p = 0.12 ; x = 4
--------------------
P(x=4) = 5C4(0.12)^4(0.88)^1 = 5*(0.00030736)(0.88) = 0.000912384..
===================
Cheers,
Stan H.